### v1

# Low-hallucination Synthetic Captions for Large-Scale Vision-Language Model Pre-training

Abstract
In recent years, the field of vision-language model pre-training has experienced rapid advancements, driven primarily by the continuous enhancement of textual capabilities in large language models. However, existing training paradigms for multimodal large language models heavily rely on high-quality image-text pairs. As models and data scales grow exponentially, the availability of such meticulously curated data has become increasingly scarce and saturated, thereby severely limiting further advancements in this domain. This study investigates scalable caption generation techniques for vision-language model pre-training and demonstrates that large-scale low-hallucination synthetic captions can serve dual purposes: 1) acting as a viable alternative to real-world data for pre-training paradigms and 2) achieving superior performance enhancement when integrated into vision-language models through empirical validation. This paper presents three key contributions: 1) a novel pipeline for generating high-quality, low-hallucination, and knowledge-rich synthetic captions. Our continuous DPO methodology yields remarkable results in reducing hallucinations. Specifically, the non-hallucination caption rate on a held-out test set increases from $48.2\%$ to $77.9\%$ for a 7B-size model. 2) Comprehensive empirical validation reveals that our synthetic captions confer superior pre-training advantages over their counterparts. Across 15 vision language tasks, the model trained with our data achieves a significant performance gain of at least $6.2\%$ compared to alt-text pairs and other previous work. In 20 common cognitive domains, the model trained with our data outperforms other counterparts by at least $6.8\%$. Meanwhile, it also offers considerable support in the text-to-image domain. With the support of our dataset, the FID score is reduced by $17.1$ on a real-world validation benchmark and $13.3$ on the MSCOCO validation benchmark. 3) We will release Hunyuan-Recap100M, a low-hallucination and knowledge-intensive synthetic caption dataset, which will enable breakthroughs in vision language pre-training, cross-modal generation, and related applications.
1 Introduction
With the enormous power of large language models, remarkable performance gains have recently been achieved in a variety of tasks in natural language processing (NLP), computer vision (CV), and also in cross-modal fields (Brown et al., 2020; Chung et al., 2022; Chowdhery et al., 2022; Touvron et al., 2023a; Dosovitskiy et al., 2021; Alayrac et al., 2022; Liu et al., 2023; Achiam et al., 2023; Wang et al., 2024b; Hurst et al., 2024). The Large Language Model (LLM) and the Vision Language Model (VLM) are usually equipped with Transformer (Vaswani et al., 2017) as the backbone and then pre-trained with a tremendous amount of unlabeled data. The strong representation ability of the model, the massive amount of data, and the effective means of training make the foundation models powerful for successfully solving the tasks of vision and language.
While text data has experienced exponential growth and fuels large language models, vision language model pre-training confronts a unique data bottleneck: high-quality text-image alignments remain scarce despite their critical role in training vision language models. Most traditional approaches to building large-scale multimodal datasets mainly rely on Web-based alt-text extraction (Sharma et al., 2018; Changpinyo et al., 2021; Ordonez et al., 2011; Schuhmann et al., 2021, 2022; Byeon et al., 2022; Gadre et al., 2023; Liu et al., 2022; Gu et al., 2022). The web-based alt-text data typically undergoes a lengthy processing pipeline, involving stages such as text quality cleaning, image low-quality filtering, and cross-modal relevance validation, ultimately producing image-text pairs. Although significant data loss occurs at each processing stage and even with these efforts, the final output still suffers from inherent limitations including a low text informativeness density and partial relevance deficiencies between visual content and the corresponding descriptions. As the naturally existing image-text pair data is exhausted, and the quality of data produced by traditional image-text extraction methods drops significantly as the scale expands, the marginal effect they generate in the training of VLMs greatly diminishes. Therefore, new methods are needed to synthesize large-scale, high-quality multimodal data.
Recently, several exploratory investigations have been carried out in multi-modal synthetic data, which have, to some extent, demonstrated their positive impact on various cross-modal tasks (Li et al., 2023a; Betker et al., 2023; Yu et al., 2024; Li et al., 2024b). However, there are significant defects in all of them, mainly reflected in the presence of many hallucinations or a relatively low level of information content. As shown in Figure 1, there are a large number of hallucinations (labeled in red) in the previous synthetic captions, and the captions are rather monotonous, lacking informative expressions. Therefore, in this paper, we focus on addressing two issues: 1. How to establish an effective pipeline for generating low-hallucination synthetic captions? 2. How can more valuable information be injected into synthetic captions to increase the information density of synthetic data?
Figure 1: Examples of the original caption, previous synthetic caption and our synthetic caption.
How to establish an effective pipeline for generating low-hallucination synthetic captions? Large-scale synthetic captions inherently depend on robust tools, and vision language models emerge as a natural and fitting choice. However, hallucination is an inherent problem in the multimodal large language model (Bai et al., 2024). As shown in Figure 1, the data in the Recap-DataComp-1B (Li et al., 2024b) dataset contains a significant amount of hallucinatory information. We further demonstrate the harmfulness of a large amount of hallucinatory data in Section 4. The inevitable and most crucial issue in caption synthesis lies in resolving the hallucination problem. Existing work (Bai et al., 2024) has discussed the causes of hallucinations from the perspectives of data, model, training, and inference. During generation, as the length of the sequence grows, self-attention will focus more on previously generated text tokens, i.e., the attention to visual content is diluted(Wang et al., 2023b, a, 2024a; Huang et al., 2024). We agree that hallucinations in VLMs are mainly caused by uncertainty in decoding visual-related intentions. Therefore, we propose the use of Direct Preference Optimization (DPO) method to eliminate hallucinations in synthesized captions by increasing the confidences of real visual contents and suppressing the confidences of hallucinatory visual contents. Fortunately, this method has proven to be highly effective. Furthermore, in the synthetic caption task, we reveal that DPO adheres to a significant scaling law within a specific data size range. Upon saturation, resampling can further elevate the quality of synthetic captions. As presented in Table 1, when applied to the same dataset, our method significantly enhances the non-hallucination rate, elevating it from 29.7% to 73.3%, in contrast to the approach used in Recap-DataComp-1B (Li et al., 2024b). We elaborate on our hallucination evaluation metrics in Section 4.1 and the continuous DPO method in Section 3.
Dataset # Length # Details $\uparrow$ # None Hallucinations $\uparrow$ # Hallucination Rate $\downarrow$
Capfusion-120M 23.5 2.58 72.3% 14%
Our Method 80.4 6.9 73.9% 5.4%
Recap-DataComp-1B 49.41 6.95 29.7% 24.9%
Our Method 72.60 6.64 73.3% 5.64%
Hunyuan-Recap100M 103.15 8.1 77.9% 4.2%
Table 1: A comparison between existing synthetic caption datasets and our proposed dataset in terms of the number of hallucinations and the hallucination rate. Here, "Length" represents the average length of the caption, "Details" represents the average number of visual details that appear in each caption, "None Hallucinations" indicates the proportion of non-hallucinatory data in the overall data, and "hallucination rate" represents the proportion of hallucinations among all details. A total of 1,000 data samples were taken, and our method reports the corresponding results on the respective data with our synthetic caption pipeline respectively.
How can more valuable information be injected into synthetic captions to increase the information density of synthetic data? The answer to this question can be rather straightforward: infuse as much knowledge as possible into the caption via a well-trained VLM. The injection of knowledge into synthetic captions can be derived from two primary sources. The first source is external text, such as alt-text, that is relevant to the image, while the second source stems from world knowledge encapsulated within a VLM. Usually, a VLM model tends to provide captions that are vague but correct. Therefore, we need to train a VLM model so that it can fully integrate the knowledge in text and its own world knowledge, ultimately generating synthetic captions with high informational value. Specifically, in our methodology, we employ a powerful VLM, namely GPT-4o, in conjunction with a meticulously crafted prompt that encompasses detailed requirements. This combination is utilized to generate labeled data, with a particular emphasis on extracting and focusing on the key information embedded within alt-text. For instance, we focus more on the name of a specific entity, using oxalic acid as an illustration here. After that, we perform manually post-processing on obvious erroneous data to ensure that the labeled data is accurate, contains the key information, and excludes the noise in the injected information. Finally, we train an open source VLM model, specifically Qwen2-VL-7B, using the training data. For details of the prompt and data processing, see Section 3.
By comprehensively leveraging the aforementioned methods, we train a recaption model and utilize it to generate 100 million pieces of data. Our model has achieved a notable enhancement in the evaluation metrics for low-hallucination quantification. Moreover, we carried out a verification of the data’s validity. The results show that vision-language models pre-trained with Hunyuan-Recap100M exhibit substantially improved overall performance compared to those using alt-text data or conventional synthetic captions, particularly showing enhanced capabilities in perception tasks.
Our contributions are as follows.
(1) We tackle the challenge of constructing a synthetic caption pipeline capable of generating captions that are both low in hallucination and rich in knowledge, thereby alleviating the scarcity of high-quality multimodal data essential for vision-language model pre-training.
(2) We propose a series of effective methods, knowledge-enriching SFT and continuous DPO, to eliminate hallucinations and inject knowledge into synthetic captions. Eventually, we trained a recaption model, whose proportion of hallucination-free results far exceeds that of the existing best models. The proportion of hallucination-free results has increased from $48.2\%$ to $77.9\%$. It also exhibits favorable performance in knowledge diversity, as reflected in a more accurate cognition of real-world entities.
(3) We will release a high-quality synthetic caption dataset comprising $100M$ image-text pairs, Hunyuan-Recap100M, with the aim of fostering further advancements within the research community in areas like vision language model pre-training, text-to-image, and text-to-video generation.
2 Related Work
Vision-Language Model. Vision Language Models (VLMs) can be classified into two distinct categories. The first category consists of models that rely on dual encoders, specifically a vision encoder and a text encoder. These models are typically trained in a contrastive fashion and exhibit remarkable proficiency in tasks such as image recognition. Prominent examples within this group include CLIP (Radford et al., 2021), FLIP (Li et al., 2023b) and SigLIP (Zhai et al., 2023b). The second category of methods focuses on aligning the visual modality with a Large Language Model (LLM). With the rapid development of LLM, the astonishing success of GPT (Brown et al., 2020; OpenAI, 2023), LLaMA (Touvron et al., 2023a, b; Grattafiori et al., 2024), and DeepSeek-LLM (Liu et al., 2024; Guo et al., 2025) has made the empowerment of large language models with visual perception an important and popular research direction. Flamingo (Alayrac et al., 2022) is the first large-scale vision language model, which aligns the visual encoder with a large language model. LLaVA (Liu et al., 2023) presents an instruction tuning methodology for training large multimodal models, thus charting the course for VLM training. Subsequent efforts have been made to refine the details of each module and continuously optimize the overall performance of the model (Wang et al., 2024b; Wu et al., 2024; McKinzie et al., 2024; Chen et al., 2024b; Liu et al., 2024). Although, compared with former multimodal models (Zeng et al., 2021; Singh et al., 2021; Diao et al., 2022; Wang et al., 2022; Zhang et al., 2023a), current LLM based vision language models are much more powerful in terms of capabilities, there has not been much improvement in data utilization; they still rely on large-scale image-text pairs. Similarly, VLMs with dual encoders, as referred to in the first category above, also require large-scale image-text pair data. However, the data of the image-text pairs are gradually being exhausted (Villalobos et al., 2022). The total amount of a large number of cleaned image-text pairs is fixed, and the expansion of data scale can hardly keep up with the requirements of model scaling. Therefore, despite the remarkable advancements achieved in the VLM training paradigm, the utilization of large-scale image-text pairs data has encountered certain limitations. To overcome this issue, a potential solution lies in synthetic data.
Synthetic data. Synthetic data has been extensively utilized in large language models, yielding remarkable results (Zhang et al., 2024; Shao et al., 2024; Zhu et al., 2024; Yang et al., 2024). In addition, there is a growing presence of synthetic data within the multimodal domain. BLIP-2 (Li et al., 2023a) employs an image caption model to synthesize short captions to serve as substitutes for alt-text, and CapFusion (Yu et al., 2024) further enhances the quality of such data using large language models. However, the captions generated through this type of work tend to be rather simplistic, merely representing minor refinements of the original alt-text without adding substantial information value. Recap-DataComp-1B (Li et al., 2024b) aims to transform alt-text using vision language models, yet it encounters the issue that existing VLMs can produce severe hallucinations. Other works focused on dense captioning, such as Allava (Chen et al., 2024a) and Dense-Fusion-1M (Li et al., 2024c), are incapable of injecting rich knowledge in a targeted fashion and, on this basis, maintain relatively low levels of hallucination. In summary, knowledge injection and elimination of hallucinations are two key challenges in the realm of multimodal synthetic data. A significant amount of research has been dedicated to exploring the hallucination problem in VLMs, including aspects such as the causes underlying hallucinations and strategies for their elimination (Bai et al., 2024). However, a comprehensive method that can simultaneously ensure content diversity and maximally eliminate hallucinations in the synthetic visual caption task remains elusive.
Our approach has achieved substantial advances in both hallucination elimination and knowledge injection. Ultimately, its efficacy has been demonstrated in the training of VLMs, establishing our caption dataset as a viable alternative for the large-scale training of vision language models. This eliminates the constraints imposed by the data scale and avoids the challenges associated with data quality degradation resulting from the expansion of the data scale.
3 Methodology
3.1 Continuous DPO
Reinforcement learning has been reported to be an effective method of eliminating hallucinations in Large Language Models (LLMs) (Zhang et al., 2023b; Huang et al., 2025). Direct Preference Optimization (DPO) (Rafailov et al., 2023) emerges as a reinforcement learning based fine-tuning paradigm for LLM that directly optimizes preference data, effectively mitigating harmful content generation, such as hallucinations. Consequently, the application of DPO to mitigate hallucinations in image captions appears to be a natural and logical approach. The optimization objectives of the DPO algorithm are shown in Equation 1 and Equation 2. As noted in Rafailov et al. (2023), the loss function gradient $\mathcal{L}_{\text{DPO}}$ increases the probability of preferred completions $y_{w}$ while simultaneously reducing the probability of dispreferred completions $y_{l}$. Intuitively, by maximizing the probability of samples free of hallucinations and minimizing the probability of samples containing hallucinations, our recaption model can generate captions that are as free from uncertainty and potential hallucinations as possible.
$\mathcal{L}_{\text{DPO}}=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{ref}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\right]$ (1)
$\nabla_{\theta}\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=$ (2)
$-\beta\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\underbrace{\sigma(\hat{r}_{\theta}(x,y_{l})-\hat{r}_{\theta}(x,y_{w}))}_{\text{higher weight when reward estimate is wrong}}\left[\underbrace{\nabla_{\theta}\log\pi(y_{w}\mid x)}_{\text{increase likelihood of }y_{w}}-\underbrace{\nabla_{\theta}\log\pi(y_{l}\mid x)}_{\text{decrease likelihood of }y_{l}}\right]\right],$
where $\hat{r}_{\theta}(x,y)=\beta\log\frac{\pi_{\theta}(y\mid x)}{\pi_{\text{ref}}(y\mid x)}$ is the reward implicitly defined by the language model $\pi_{\theta}$ and the reference model $\pi_{\text{ref}}$.
However, in our task, the DPO algorithm does not demonstrate continuous performance improvement with respect to scale. Specifically, after initial rapid scaling up of performance, its performance plateaus, as illustrated in Figure 4. We turn our attention back to Equation2. Here, the examples are weighted according to how much higher the implicit reward model $\hat{r}_{\theta}(x,y)$ rates the dispreferred completions. As we continuously expand the scale of the DPO data, the implicit reward $\hat{r}_{\theta}(x,y)$, which serves as the optimization weight, does not change significantly. Consequently, after reaching a local optimum, the model ceases to optimize further. Therefore, we make slight modifications to the DPO algorithm and propose Continuous DPO (CDPO). Once the performance reaches a bottleneck phase, we switch to a different reference model, resample data using the current model, adjust the implicit reward model $\hat{r}_{\theta}(x,y)$ accordingly, and then re-optimize $\pi_{\theta}$. Furthermore, when CDPO training with preference data is performed, in order to prevent the sample length from serving as a learning signal (e.g., where longer captions may contain more hallucinations), we conduct an additional sequence-length balancing operation on the samples. This ensures that the average lengths of the positive and negative samples are close to each other.
3.2 Knowledge-enriching SFT
Previous recaption models (e.g.(Li et al., 2023a; Yu et al., 2024)) often produce generalized and imprecise descriptions, such as "a man performing on stage" in Figure 1. This type of descriptive information has limited informational value and offers little assistance in training vision language models. Hence, in our work, we place significant emphasis on the integration of valuable information. We believe that valuable information is mainly reflected in two aspects: one is the richness of descriptive details and the other is the knowledge content of the description. The former is easy to understand, while the latter is manifested in the supplementation of external knowledge related to the image. For example, in Figure 1, in the case of "oxalic acid", we hope to describe it as follows: The image shows the molecular chemical structure of oxalic acid, with the chemical formula $C_{2}H_{2}O_{4}$. The oxalic acid molecule consists of two carbon atoms, two hydrogen atoms, and four oxygen atoms. In the structural formula, the two carbon atoms are connected by single bonds, and each carbon atom is attached to a hydroxyl group ($OH$) and a carbonyl group ($C=O$), which are the typical structural features of oxalic acid.
Figure 2: A GPT-4o prompt designed to construct image captions rich in descriptive details. In the prompt, the red-highlighted portion lays stress on the methodology of knowledge injection. Furthermore, it guarantees that the ultimately generated caption incorporates a rich body of world knowledge via a two-step procedure.
By constructing SFT data, captions rich in descriptive details can be generated, enabling the model to learn to produce longer captions that contain as much specific information as possible through fine-tuning. Therefore, we carefully constructed a complex prompt and obtained training data using GPT-4o, see Section 3.3. Note that although GPT-4o can understand the requirements of the prompt, the output data still contain a substantial amount of hallucinations. As a result, we perform a manual review. Moreover, to incorporate knowledge into captions more effectively, we screened the original image-text pairs. In addition to general alt-text, we selected object-related data with a higher informational value [^1] In our study, we primarily regulate this aspect through the meticulous selection of alt-text data sources. For example, we deliberately choose a diverse range of encyclopedia data. This allows the descriptions in the captions to focus more on specific object descriptions, such as the "tragocephala pretiosa" in Figure 1. After reinforcement of the above two aspects, we construct knowledge-enhanced supervised fine-tuning data capable of supporting the training of image captions with rich content.
3.3 Training Pipeline
As mentioned above, our comprehensive training pipeline consists of two sequential stages: knowledge-enriching SFT followed by continuous DPO. In this section, we introduce these stages specifically. The general pipeline of our work is shown in Figure 3. We adopt Qwen2-VL-7B (Wang et al., 2024b) as the base model and use the LLaMA Factory (Zheng et al., 2024) platform to carry out the overall training process.
Figure 3: The illustration of our recaptioning pipeline
Knowledge-enriching SFT. We sampled raw image-text pairs from open-source multimodal datasets including CC3M (Sharma et al., 2018), CC12M (Changpinyo et al., 2021), Datacamp (Gadre et al., 2023) Wukong (Gu et al., 2022), and Wikipeda (Srinivasan et al., 2021). Using the prompt shown in Figure 2, we employed GPT-4o to extract the image-alttext-caption data. Following a meticulous manual review process, $43,408$ data entries were retained. Subsequently, we performed LoRA fine-tuning on the Qwen2-VL-7B model. Specifically, the learning rate was set at $51.0e-$ 5, the global batch size was configured as $128$, and the number of training epochs was established at $10$. Through this process, we eventually acquired our Recaption-SFT model.
Continuous DPO. As outlined in Section 3.1, our methodology begins with training a DPO intermediate model. We first randomly sample $300,000$ image-text pairs and then employ the Recaption-SFT model for inference. To induce controlled stochasticity, we systematically modulate three sampling hyperparameters, TOPP=$1.0$, TOPK=$20$, and TEMPERATURE=$1.0$. Each image-text pair undergoes eight parallel inference iterations, thereby generating eight divergent candidates. We construct preference-ordered pairs by selecting the best and worst outputs from each set with a critic model  [^2] It is an in-house expert system that primarily makes point-by-point judgments by dissecting the details in captions and then assesses the hallucination level of the details. The main pipeline is similar to that of Section 4.1, but it does not use GPT-4o.
To eliminate length-induced training bias, we implement length-balanced sampling that retains $218k$ high-quality pairs, enforcing alignment of length distribution between positive and negative samples. Ultimately, the curated dataset was used for DPO training on the Recaption-SFT model, resulting in the DPO intermediate model.
Upon acquiring the intermediate model, we further sampled $200k$ data instances from the data sources and used this intermediate model to generate preference data with the same hyperparameters and the judge pipeline. Following the length-balancing strategy, we ultimately retained $139k$ data instances for continuous DPO training. All hyperparameters at different stages are listed in Table 2.
Stage-1-SFT Stage-2-DPO Stage-2.5-CDPO
Resolution # pixels [3136, 12845056] [3136, 12845056] [3136, 12845056]
Dataset # Samples 43K 218K 139K
Finetuning Type LoRA LoRA LoRA
Batch Size 128 64 64
Learning Rate $51e-$ 5 $65e-$ 6 $65e-$ 6
Learning Epoch 10 1 1
Table 2: Detailed configuration for each training stage of our recaption model. The table outlines the progression of vision parameters, dataset characteristics, model specifications and training hyperparameters.
4 Experiments
This section outlines our quality assessment criteria for recaption data and elaborates on our work from two aspects: the intrinsic quality of the dataset and its effectiveness in large-scale vision language model pre-training.
(a) Non-hallucination Rate
(b) Low-hallucination Rate
Figure 4: Performance of different DPO strategies. The horizontal axis represents the scale of the training data, and the vertical axis represents the proportion of hallucination-free captions on the validation set (in percentage). The CDPO strategy uses the same training data as DPO in the early stage of training and resamples $139k$ new data after training $218k$ data.
4.1 Evaluation Metrics
We believe that the quality measurement of recaption data encompasses two aspects. One is the assessment of harmful information, namely hallucinations. This information should be minimized as much as possible. The other aspect is the measurement of content richness, which can usually be verified by the effectiveness of model training. Therefore, in terms of the quality of the recaption data itself, we pay more attention to the indicators of hallucinations, while the measurement of its richness will be reflected through the training effect of large-scale vision language models in Section 4.3. We take advantage of the hallucination assessment method of CIEM (Hu et al., 2023) and leverage the existing large language model, GPT-4o, to decompose potential detailed elements within the caption. Subsequently, we use GPT-4o to make judgments about the presence of hallucinations. Finally, we count the number of hallucinations in a caption. We define the proportion of captions that do not contain hallucinatory details as the non-hallucination rate, and the proportion of captions that contain at most two hallucinatory details as the low-hallucination rate.
4.2 Performance of Recaption Model
We first report the evaluation results of our recaption model on open-source data, which come mainly from Wukong (Gu et al., 2022). As illustrated in Figure 4, the base model (Qwen2-VL-7B) exhibits a considerably severe level of hallucination. Specifically, its non-hallucination rate ($48.33\%$) and low-hallucination rate ($87.87\%$) are quite low. Following Supervised Fine-Tuning, the low-hallucination rate increases significantly. Subsequently, with the application of DPO, both the non-hallucination rate and the low-hallucination rate are substantially improved, by $34.03\%$ and $10.21\%$, respectively. Furthermore, we observe that DPO demonstrates a notable scaling law during the initial stage of the recaption task. However, the performance plateaus once the data size attains a certain scale. Our CDPO algorithm can further enhance both the non-hallucination rate (by $+3.4\%$) and the low-hallucination rate (by $+0.5\%$).
Meanwhile, we also use our model to infer on the same image-text pairs as other recaption works, e.g. Capfusion (Yu et al., 2024) and Recap-DataComp-1B (Li et al., 2024b), and report the results in Table 1. For the same original alt-text data, the recaption data generated by our model contain significantly more details than that of Capfusion and have much fewer hallucinations than Recap-DataComp-1B. Specifically, when dealing with the same data, compared to Capfusion, the number of details in the captions of our model increases by $267\%$, the average length of our captions is $3.5$ times longer than that of Capfusion, and the rate of hallucinatory details decreases by $8.6\%$. Moreover, compared to Recap-DataComp-1B, the non-hallucination rate increases by $43.6\%$. Note that since the length of captions in Capfusion is short, its non-hallucination rate is quite high. However, because of the lack of details, its value is limited.
Scientific QA OCR QA Benchmark QA Hallucination Tasks
Dataset MathVista MMMU AI2D ScienceQA OCRBench DocVQA TextVQA InfographicsVQA ChartQA MMBench MMStar RealworldQA MMVet HallusionBench None-Hallu. Rate Avg.
testmini val test test test val val val test (aug./hum.) en-dev test test test test
Alt-text(20M) 42.20 42.24 67.06 81.75 49.0 65.01 55.97 31.17 46.92 67.96 44.27 57.12 43.81 34.29 68.38 53.12
Recap-DataComp-1B(20M) 51.40 43.33 68.60 86.27 55.8 46.90 64.75 33.52 53.16 66.58 48.46 58.95 39.66 29.84 33.79 52.07
Hunyuan-Recap100M(20M) 52.50 44.27 73.59 85.82 57.3 66.27 72.59 42.47 57.00 72.77 50.13 60.26 46.02 35.45 73.87 59.35
Hunyuan-Recap100M 45.10 43.56 73.19 85.92 65.80 66.30 72.42 44.66 56.0 74.83 49.13 64.44 49.00 34.60 72.61 59.83
Alt-text(20M)$\ast$ 50.10 45.0 73.18 85.57 59.8 77.34 68.47 43.99 62.08 72.42 47.89 61.57 42.09 33.80 39.27 57.50
Recap-DataComp-1B(20M)$\ast$ 53.70 44.44 74.65 86.12 62.1 79.11 66.39 45.38 64.28 65.89 50.44 61.96 39.34 32.77 43.62 58.01
Hunyuan-Recap100M(20M)$\ast$ 56.20 44.00 77.91 89.74 65.9 83.06 71.37 50.32 68.96 72.34 54.66 65.49 43.17 36.38 71.86 63.42
[…]
Table 3: Experimental results on 15 vision-language tasks. Avg. is the average score of all vision-language tasks. The None-Hallu. Rate is a hallucination assessment metric introduced in Section 4.1. We randomly sampled 1000 pieces of DataCamp data to evaluate the level of hallucination. The text in bold represents the optimal performance for the task under the 20M setting, whereas the text with an underline signifies further enhancement when the entire dataset is utilized. The $\ast$ indicates that a unified lite supervised fine-tuning (SFT) (Wang et al., 2023c; Li et al., 2024a) has been performed on the basis of pre-training. Note that our main purpose is to verify the impact of different caption datasets on the training of VLM models. Therefore, the above results are comparative experiments conducted under a fixed data setting (20M) without carefully supervised fine-tuning and reinforcement learning and are not comparable to SOTA models.
4.3 Performance of Recaption Data
We validate the effectiveness of Hunyuan-Recap100M in two scenarios, vision-language model pre-training and text-to-image generation. This validation is conducted primarily from the viewpoints of model training efficacy, hallucination levels, and knowledge coverage.
Vision-Language Model Pre-Training. We first validate the effectiveness of our data set in the pre-training of vision language models. The aim is to demonstrate that, compared to alt-text and other recaption data available on the market, Hunyuan-Recap100M can better assist in the training of vision language models and achieve better performance in both general open source multimodal tasks and in-house knowledge-related test sets. In the VLM framework, we select the LLaVA (Liu et al., 2023) architecture, with SigLIP (Zhai et al., 2023a) as visual encoder and Hunyuan-7B as the base large language model. To make a fair comparison of the impact of caption data, we sample $20M$ data points from alt-text/recaption data. Note that the alt-text group experiment uses the same images as Hunyuan-Recap100M, while Recap-Datacamp-1B uses randomly sampled data [^3] Since our data has undergone knowledge screening, its distribution is not exactly the same as that of DataCamp. As shown in Table 3, on 15 open source multimodal tasks, the VLM model trained with Hunyuan-Recap100M significantly outperforms those trained with alt-text or Recap-Datacamp-1B. Especially in hallucination-related tasks, Hunyuan-Recap100M shows a substantial lead over Recap-Datacamp-1B. Additionally, we evaluated the model’s convergence behavior through loss function monitoring. When using identical supervised fine-tuning datasets, models pre-trained with Hunyuan-Recap100M demonstrated markedly improved convergence properties.
In addition, to further measure the improvement in the perceptive level of the VLM caused by Hunyuan-Recap100M, we manually collected pictures with different visual characteristics from 20 domains, including animals, plants, clothing, landmarks, etc., totaling 1,197 pieces of data with manually annotated labels. We conducted an evaluation of the perceptive ability of the VLM in this set of data and the results are shown in Table 4. The VLM model trained with Hunyuan-Recap100M significantly outperforms other counterparts in 18 of the 20 categories. In general, it brings at least a $6.8\%$ improvement on average. After fully applying Hunyuan-Recap100M, the model has shown significant improvements in 17 of 20 perception tasks, with an additional average improvement of $9\%$.
Dataset Life Landmark Celebrity Food Cloth Good Game Traffic Art Festival Moive Icon Animal Plant Vegetables Sport Medical Military Tool History Avg.
Alt-text(20M) 34.1 35.9 23.4 22.4 26.2 22.7 24.5 34.5 23.3 27.1 25.5 27.5 54.2 27.3 22.2 65.5 32.6 29.4 20 27.3 30.0
Recap-DataComp-1B(20M) 13.6 19.7 18.9 4.5 29.9 23.2 21.6 33.9 12.2 10.0 12.7 17.0 49.2 13.6 13.1 57.9 19.6 19.6 30.4 17.6 21.8
Hunyuan-Recap100M(20M) 25.0 38.1 26.3 28.4 35.4 41.2 36.5 62.5 25.0 18.5 33.3 40.4 60.7 31.8 31.7 72.4 27.3 34.0 36.5 31.4 36.8
Hunyuan-Recap100M 40.0 44.3 39.0 34.4 53.8 38.1 51.9 63.6 27.5 35.1 31.5 51.0 66.1 50.7 50.8 79.3 46.3 39.6 36.4 38.2 45.8
Table 4: Experimental results on identification of 20 common vision object categories. The text in bold represents the optimal performance for the task under the 20M setting, whereas the text with an underline signifies further enhancement when the entire dataset is utilized.
Text-to-Image Model Training We also verify the effectiveness of Hunyuan-Recap100M on the text-to-image generation task. We perform fine-tuning on the technical stack of Hunyuan-DiT (Li et al., 2024d) in order to demonstrate that our data can help generate specific visual images. As evidenced in Table 5, the publicly released Hunyuan-DiT model can significantly achieve a lower FID and a higher CLIP score on the MSCOCO dataset, as well as our internal perceptual test, when using our recaption data. Furthermore, after fine-tuning the data with Hunyuan-Recap100M, there is a further improvement in both the FID and the CLIP Score. This finding indicates that our data can offer considerable assistance in the realm of world perception and also enable text-to-image models to achieve better performance in terms of authenticity. We present realistic cases in Section B.
[…]
Dataset Row Caption Our Recaption Fine-tuning with Our Recaption Row Caption Our Recaption Fine-tuning with Our Recaption
perceptive Test Set 62.38 49.13 45.33 0.323 0.344 0.357
MSCOCO Test set 28.79 21.09 15.46 0.307 0.305 0.313
Table 5: Text-to-Image Evaluation of Hunyuan-DiT on our internal perceptual test and MSCOCO-30K, w/wo fine-tuned with our recaption data. Note that, due to computational resource limitations, we only sampled $2M$ data from Hunyuan-Recap100M for LoRA fine-tuning.
5 Conclusion and Limitation
5.1 Conclusion
In this work, we tackle the issue of synthesizing captions that are rich in knowledge and exhibit low levels of hallucination. We present a novel pipeline for high-quality synthetic captions, featuring two effective techniques: continuous DPO and knowledge-enriching SFT. We elucidate the beneficial role of DPO in alleviating hallucinations within image captions, along with its scaling law within a certain scope. Furthermore, we propose an approach to continually improve model performance once the scaling law ceases to be effective. Experimental results demonstrate that Hunyuan-Recap100M outperforms original alt-text and other synthetic captions in the pre-training of vision language models by a large margin. We will make Hunyuan-Recap100M publicly available to promote the development of multimodal related fields in the open source community.
Limitations.
Like most of the existing work on vision language models, the entire project consumed a lot with high electricity costs, although we only tested a $7B$ size model on $100M$ training data. Due to limited computing power, we did not scale up the use of recaption data to a large extent in this paper. We leave this work for future research. We also did not verify the effectiveness of the large-scale recaption data we generated for training text-to-image models from scratch. This will also be considered as future work.
Potential Risks.
The original image-text pairs are mainly derived from open source datasets. Although we have taken measures to detoxify to a great extent, there are still potential risks, especially in the field of image generation, such as fake portraits on social media (Hill and White, 2020), which is a common potential risk in image generation research.
Appendix A Loss of Vision-Language Model Training
To address the issue of incomparability of loss arising from diverse data distributions, we examine the convergence of pre-trained models that are trained on different datasets but are evaluated on the same supervised fine-tuning data, as illustrated in Figure 5. In particular, it can be observed that models trained with Hunyuan-Recap100M display a better convergence performance when dealing with the same data from downstream tasks. Specifically, these models exhibit lower loss values at both the starting point, with a loss value of $0.925$, and the convergence point, where the loss value is $0.48$.
Figure 5: Loss curve of vision-language model pre-trained with alt-text, Recap-DataComp-1B or Hunyuan-Recap100M. Note that all pre-trained models use the same amount of data (20M), and at the same time, the three curves are evaluated on exactly the same supervised fine-tuning data.
Appendix B Examples of Text-to-Image Generations
In this section, we present a series of examples of text-to-image generation in Figure 6. The synthetic captions corresponding to the text concepts in Figure 6 can be found in Table 6. From these examples, we can see that the HunyuanDiT model does not respond to a few concepts, such as abacus and Yueqin, and its depictions of some other concepts are also inaccurate, like jellyfish and acerola. After tuning with Hunyuan-Recap100M, the model’s understanding of these concepts becomes more precise, and its generated image style is also more realistic. The situations demonstrated in these cases are consistent with the reduction in FID shown in Table 5, highlighting the wide coverage of concepts perceived by our data. This is corroborated by the significant improvement in the performance of the VLM trained with Hunyuan-Recap100M on numerous object perception tasks, as presented in Table 4.
Text Image HunyuanDiT HunyuanDiT wRecap HunyuanDiT wFT
Jellyfish Refer to caption Refer to caption Refer to caption Refer to caption
Acerola Refer to caption Refer to caption Refer to caption Refer to caption
Abacus Refer to caption Refer to caption Refer to caption Refer to caption
Yueqin Refer to caption Refer to caption Refer to caption Refer to caption
Alpaca Refer to caption Refer to caption Refer to caption Refer to caption
The Grotto Refer to caption Refer to caption Refer to caption Refer to caption
Shanghai Tower Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Demonstration of text-to-image cases. Here, the original instruction for image generation is denoted as "Text", the original picture as "Image", the generation effect of the original open-source model as "HunyuanDiT", the generation effect of the original open-source model with our Synthetic texts as "HunyuanDiT wRecap", and the effect of the model fine-tuned with our recaption data as "HunyuanDiT wFT".
Text Our Recaption
Jellyfish The picture shows two jellyfishes. Their bodies are orange or light red with distinct stripes or lines. Their tentacles are long, slender, transparent or semi-transparent. The background is black, highlighting the colors of the jellyfish and the details of their tentacles. The jellyfish are swimming or floating in the water, and their tentacles are spread out in the water, forming a distinct linear effect.
Brazilian Acerola The image shows the branches of a tree laden with red and green fruits. The red fruits look ripe, while the green ones appear unripe. There are also green leaves on the branches. In the background, more green leaves and a few blurred flowers or small buds can be seen. The overall environment seems to be an outdoor natural setting.
Abacus There is a group of people sitting in front of a wooden table in the picture, using abacuses for learning or operation. There are several abacuses on the table. Wooden doors and windows can be seen in the background, indicating that this is an indoor environment, perhaps a classroom or a study place. People are wearing casual or semi-formal clothes, looking relaxed.
Yueqin This is a Yueqin, which has a circular resonator and four strings. There are multiple positions for pressing the strings on the body of the instrument, and there are four tuning pegs on the headstock. Overall, it exhibits the characteristics of a traditional musical instrument.
Alpaca There are two alpacas standing on the grass in the image, with blue sky and white clouds in the background. The alpaca on the left has lighter fur, while the one on the right has darker fur. The grass is green, and the postures of the alpacas suggest that they are either walking or standing.
The Grotto The Grotto is a church located in a natural environment. The facade of the church is a huge rock cave, covered with green vegetation above the cave, and surrounded by dense trees. Inside the church, there is an altar, above which a cross is hung, surrounded by lighting. In front of the church, there is a row of wooden benches for believers to sit and pray.
Shanghai Tower Shanghai Tower is a skyscraper with a spiral glass curtain wall design on its exterior. At the top, there is a circular observation deck, surrounded by other high-rise buildings. The overall architectural style is modern and futuristic.
Table 6: The original text of the image in Figure 6 and the synthetic caption after being processed by our recaption model.